# k* Distribution: Evaluating the Latent Space of Deep Neural Networks using Local Neighborhood Analysis

### v1

@@ Abstract
Most examinations of neural networks’ learned latent spaces typically employ dimensionality reduction techniques such as t-SNE or UMAP. While these methods effectively capture the overall sample distribution in the entire learned latent space, they tend to distort the structure of sample distributions within specific classes in the subset of the latent space. This distortion complicates the task of easily distinguishing classes identifiable by neural networks. In response to this challenge, we introduce the k* Distribution methodology. This approach focuses on capturing the characteristics and structure of sample distributions for individual classes within the subset of the learned latent space using local neighborhood analysis. The key concept is to facilitate easy comparison of different k* distributions, enabling analysis of how various classes are processed by the same neural network. This provides a more profound understanding of existing contemporary visualizations. Our study reveals three distinct distributions of samples within the learned latent space subset: a) Fractured, b) Overlapped, and c) Clustered. We note and demonstrate that the distribution of samples within the network’s learned latent space significantly varies depending on the class. Furthermore, we illustrate that our analysis can be applied to explore the latent space of diverse neural network architectures, various layers within neural networks, transformations applied to input samples, and the distribution of training and testing data for neural networks. We anticipate that our approach will facilitate more targeted investigations into neural networks by collectively examining the distribution of different samples within the learned latent space.
@@
Deep Learning, Convolution Neural Networks, Image Classification, Latent Space Visualization, Local Neighborhood Analysis
**footnotetext: These authors contributed equally to this work
@@ I Introduction
Figure 1: Overview of three distinct basic patterns of k* Distribution. Here we define the k* value of a sample point as ${}^{\text{th}}$-closest neighbor, which differs in class compared to the test point, i.e., the neighbor (sample) which breaks homogeneity in the local neighborhood of the test point. Pattern A (🟊normal-$\bigstar$) which has negatively skewed k* distribution (skewed towards high k* value) representing an ‘Fractured’ distribution of samples in latent space; Pattern B (normal-$\clubsuit$) which has almost uniform k* distribution representing a ‘Overlapped’ distribution of samples in latent space; Pattern C (normal-$\spadesuit$) which has positively skewed k* distribution (skewed towards low k* value) representing a ‘Clustered’ distribution of samples in latent space.
Figure 2: Overview of the framework to create k* Distribution. We use the learned features of a neural network to compute k* values of individual evaluated sample and then compute the k* distribution for a particular class.
@@
A significant portion of neural network research relies on creating tools to comprehend the acquired latent space and unveil the inner workings of neural networks, which are often viewed as black boxes. Nevertheless, if researchers are equipped with the essential tools to grasp the learned latent space, it becomes feasible to delve deeper, uncovering insights and reasons that can guide further research in the field of neural networks. Analyzing the neural network’s latent space poses challenges due to its intricate non-convex characteristics. However, directly evaluating and comparing the configuration and distribution of data within high-dimensional latent spaces persists despite the development of various visualization tools.
@@
Existing tools often rely on dimensionality reduction techniques like tSNE [1] and UMAP [2] to create $2$ or $3$-dimensional scatter plots for individual latent spaces, offering a broad overview. They rely on dimensionality reduction methods rooted in manifold learning which take into account specific manifold characteristics, preserving them while modifying other factors and attributes.
@@
Analyzing the configurations and structures of various sample distributions associated with a specific class within the latent space presents a challenge using existing visualization methods. This challenge extends to the comparison of multiple sample distributions which limits the analysis of the latent spaces. Additionally, visually comparing multiple latent spaces side by side can be overwhelming and confusing, especially when dealing with numerous points and varying degrees of distortion [3, 4, 5, 6, 7].
@@
If there is too much information kept, it is hard to understand. At the same time, if the information is filtered enough, there is always some perspective lacking. Therefore, some complementary tools might be crucial in interpreting these complex latent spaces.
@@
This article introduces such a tool designed to enhance the understanding of the distribution of samples associated with ‘classes/features’ in the learned latent space generated by neural networks. The proposed approach involves leveraging the local neighborhood relationships among these samples in the latent space. More precisely, the nearest-neighbor method is applied to the learned latent space to identify a k*-nearest neighbor (sample) within the local neighborhood, featuring a different class from the test sample. This process evaluates the disruption of the homogeneity within the local neighborhood of the tested sample. Subsequently, a distribution of k*—referred to as the k* distribution—is generated, incorporating all the samples belonging to a given class. Through an analysis of the k* distribution, three distinct patterns of distribution of samples in the learned latent space are identified as illustrated in Figure 1:
@@
Pattern A (🟊normal-$\bigstar$) representing Fractured distribution of samples in latent space,
@@
Pattern B (normal-$\clubsuit$) representing Overlapped distribution of samples in latent space, or
@@
Pattern C (normal-$\spadesuit$) representing Clustered distribution of samples in latent space.
@@ II Related Works
II-A Visualising the Latent Space of Neural Networks using Dimensionality Reduction Techniques:
Given that neural networks operate in a high-dimensional space, visualizing their latent space directly is challenging. To overcome this limitation, various researchers employ dimensionality reduction techniques to represent the latent space in 2 or 3 dimensions. There are a lot of techniques for visualizing the latent space like, [1, 2, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25].
Many of these algorithms require hyperparameter tuning for optimal visualization, making it challenging to ensure a fair comparison between latent spaces of varying dimensions. Additionally, while these visualizations effectively capture some perspective on the structure of the learned latent space, using them to draw a clear comparison between different local structures is challenging [3, 4, 5, 6, 7].
The effectiveness of dimensionality reduction techniques becomes apparent in instances where the latent space is well-organized and has successfully assimilated the intended information. In such cases, these visualizations demonstrate utility by aligning with pre-established interpretations. Conversely, when applied to latent space lacking clearly defined pre-established knowledge, the efficacy diminishes for such visualizations where the lack of known structure leads to a ‘blob of points’ in the visualization [7].
II-B Visualising Association of Features in the Latent Space of Neural Network:
@@
Additionally, interpretable concepts, defined as groups of latent variables in the space that are meaningful, may manifest in the latent space, further contributing to our understanding [26, 27, 31, 32, 33]. In the domain of natural language processing, these interpretations often provide insight into learning semantic and linguistic concepts provided by the lexical object [34, 35, 36]. These interpretations are often effectively visualized using analytic systems like [37, 38, 39, 40, 41, 42, 43], however, the challenge to draw a clear comparison between different local structures and multiple latent spaces persists.
@@
To examine the distribution of samples in the neighborhood of a specific sample and analyze the distribution of samples sharing a predefined label, we suggest employing the k* distribution. With our proposed approach, we present a framework (see Figure 2) for effectively comparing various latent spaces and sub-spaces with diverse sample distributions. This framework enables meaningful comparisons between the distribution of samples belonging to different classes and multiple latent spaces, offering additional insights into latent spaces beyond what existing visualizations provide.
III k* Distribution: Analysing Homogeneity in the Local Neighborhood of Samples
@@
We propose a methodology based on the local neighborhood to analyze the hyper-dimensional latent space learned by neural networks (see Figure 2). The distribution of samples and clusters in the learned latent space is analyzed by associating them with classes. Through this approach, we gain insights into the distribution patterns (see Figure 1) within the learned latent space and identify the clusters formed, thereby enhancing our comprehension of the latent space.
Visualizations inevitably involve trade-offs as they strive to make information comprehensible from specific perspectives. The concept of nearest neighbors is arguably the most relevant when portraying the structure of the distribution of samples, like clustering or fragmentation in a latent space. Alternative measures like the absolute distance can offer insight into other facets of the latent space, but they may not effectively reveal the intricate high-dimensional fracturing. Further projections of high dimensional latent space using these distances can obscure the crucial information about neighboring samples.
The Nearest Neighbours method, a widely recognized non-parametric technique, is employed enabling us to understand the positioning of latent variables in the vicinity of a latent variable in space and their corresponding class distribution. This method aids in gauging the relative distances between latent variables and grouping them into clusters. The underlying principle is grounded in the notion that local neighborhoods offer a reasonable estimate of the sample distribution within the latent space.
@@
By exploring the concept of neighborhoods and their utility in validating the local relationship among features in latent space, we introduce the k* distribution in this context. The k* distribution is constructed by taking a latent space as input and analyzing the neighborhood of a sample as shown in Figure 3.
The measurement of the nearest neighbor index allows us to address the sparsity inherent in high-dimensional space. Rather than relying on absolute distance values between points, a relative measure through the concept of neighborhood is utilized. This approach facilitates the comparison of two distinct latent spaces with varying dimensionalities. Importantly, the neighborhood concept is dimensionality-independent, making this technique applicable and effective across both low and high dimensions.
@@
The index of ${}^{\text{th}}$ nearest neighbor (sample) that belongs to a different class than the test sample disrupting the uniformity of the test sample’s neighborhood is quantified. This index is termed the k* value. Essentially, a high k* value indicates that the sample is surrounded by similar points that share the same class. In other words, a high k* value represents that a neighbor with a different class will be situated at a considerable distance from the neighborhood of the specified sample. In particular, the homogeneity of each class cluster within the learned latent space can be assessed using k* (see Figure 3). Additionally, one can ascertain whether a cluster remains cohesive for a given class or if it is fragmented across various spatial locations.
@@
Let us consider a collection of sample-label pairs $X$: $(x_{1},Y_{1}),(x_{2},Y_{2}),...,(x_{n},Y_{n})$ where $x$ is the input sample and $Y$ is the label for the input sample. Here, the latent space is embedded with such $x$ points. Let $S$ be the set of all samples $x_{p}\in X$ such that they have the same label $c$, i.e.,
$S_{c}=\{x_{i}\mid\forall x_{i}\in X~{}\text{such that}~{}Y_{i}=c\}$ (1)
@@
$D(x_{p})=\{\text{distance}(x_{p},x_{i})\mid\forall(x_{i},y_{i})\in X\}$ (2)
where $\text{distance}(a,b)$ is the distance between two samples $a$ and $b$. The commonly employed distance function is the Minkowski distance, which is a generalization of various distance metrics, including Euclidean and Cityblock distances. The Minkowski distance of order $p$ between two points $a=(a_{1},a_{2}\ldots a_{r})$ and $b=(b_{1},b_{2}\ldots b_{r})$ in $r$ dimensional space is given by,
$\text{distance}(a,b)_{\text{Minkowski},p}=\left(\sum_{i=1}^{n}|a_{i}-b_{i}|^{r}\right)^{1/r}$ (3)
Here, when $p=2$, it represents Euclidean distance ($l_{2}$ norm); when $p=1$, it represents Cityblock distance ($l_{1}$ norm); and when $p=\infty$, it represents Maximum Norm distance ($l_{\infty}$ norm).
$i^{th}$ neighbour sample $x_{i}^{p}$ to sample $x_{p}$ is defined as,
$x_{i}^{p}=x_{k}\quad\text{where}\quad k\in\operatorname*{arg\,min}_{x_{k}\in P_{i}}~{}\text{distance}(x_{p},x_{k})$ (4)
such $\text{such that}\quad P_{i}=X-\{x_{j}^{p}\mid\forall j<i\}$
Similarly, local neighbourhood space $N_{p}$ of sample $(x_{p})$ can be defined as,
@@
$\text{k}^{*}(S_{c})=\left\{\frac{\text{k}^{*}_{p}}{|S_{c}|}\mid\forall x_{p}\in S_{c}\right\}$ (7)
Here, $|S_{c}|$ is the cardinality of set $S_{c}$ representing the number of samples belonging to class $c$.
Based on the defined k* distribution we define certain metrics over it described below,
Mean of k* distribution
$\mu_{k^{*}}$ can be defined as,
@@
Standard Deviation of k* distribution
$\sigma_{k^{*}}$ can be defined as,
@@
measuring the amount of variation or dispersion in k* values.
Fischer-Pearson Skewness Coefficient of k* distribution
$\gamma_{k^{*}}$ can be defined as,
$\gamma_{k^{*}}=\frac{\frac{1}{|S|}\sum(\text{k}^{*}(S)-\mu_{k})^{3}}{\left(\frac{1}{|S|}\sum(\text{k}^{*}(S)-\mu_{k})^{2}\right)^{3/2}}$ (10)
@@
Figure 4: Visualization of the distribution of samples in latent space using, (Left) k* distribution, and (Right) Dimentionality Reduction techniques like t-SNE (Top Left), Isomap (Top Right), PCA (Bottom Left), and UMAP (Bottom Right)  of all classes of 16-class-ImageNet for the Logit Layer of ResNet-50 Architecture. Note, how the distribution of samples for a particular class is easier to compare using k* distribution compared to using dimensionality reduction techniques.
Pattern A (🟊normal-$\bigstar$) Fractured distribution of samples:
@@
Pattern B (normal-$\clubsuit$) Overlapped distribution of samples:
@@
Pattern C (normal-$\spadesuit$) Clustered distribution of samples:

### v2

@@ Abstract
Most examinations of neural networks’ learned latent spaces typically employ dimensionality reduction techniques such as t-SNE or UMAP. These methods distort the local neighborhood in the visualization, making it hard to distinguish the structure of a subset of samples in the latent space. In response to this challenge, we introduce the k* distribution and its corresponding visualization technique This method uses local neighborhood analysis to guarantee the preservation of the structure of sample distributions for individual classes within the subset of the learned latent space. This facilitates easy comparison of different k* distributions, enabling analysis of how various classes are processed by the same neural network. Our study reveals three distinct distributions of samples within the learned latent space subset: a) Fractured, b) Overlapped, and c) Clustered, providing a more profound understanding of existing contemporary visualizations. Experiments show that the distribution of samples within the network’s learned latent space significantly varies depending on the class. Furthermore, we illustrate that our analysis can be applied to explore the latent space of diverse neural network architectures, various layers within neural networks, transformations applied to input samples, and the distribution of training and testing data for neural networks. Thus, the k* distribution should aid in visualizing the structure inside neural networks and further foster their understanding.
@@
Neural Networks, Latent Space Visualization, Local Neighborhood Analysis, Class Representation, Cluster Analysis
@@ I Introduction
Figure 1: Overview of three distinct basic patterns of k* Distribution. Here, we define the k* value of a sample point as the kth-closest neighbor, which differs in class compared to the test point, i.e., the neighbor (sample) which breaks homogeneity in the local neighborhood of the test point. Pattern A (🟊$\bigstar$) which has positively skewed k* distribution (skewed towards low k* value) representing an ‘Fractured’ distribution of samples in latent space; Pattern B ($\clubsuit$) which has almost uniform k* distribution representing a ‘Overlapped’ distribution of samples in latent space; Pattern C ($\spadesuit$) which has negatively skewed k* distribution (skewed towards high k* value) representing a ‘Clustered’ distribution of samples in latent space.
@@
Asignificant portion of neural network research relies on creating tools to comprehend the acquired latent space and unveil the inner workings of neural networks, often viewed as black boxes. Nevertheless, if researchers are equipped with the essential tools to grasp the learned latent space, it becomes feasible to delve deeper, uncovering insights and reasons that can guide further research in neural networks. Analyzing the neural network’s latent space poses challenges due to its intricate non-convex characteristics. However, directly evaluating and comparing the configuration and distribution of data within high-dimensional latent spaces persists despite the development of various visualization tools.
@@
Existing tools often rely on dimensionality reduction techniques like t-SNE [1] and UMAP [2] to create $2$ or $3$-dimensional scatter plots for individual latent spaces, offering a broad overview. They rely on dimensionality reduction methods rooted in manifold learning, which considers specific manifold characteristics and preserves them while modifying other factors and attributes.
@@
Analyzing the configurations and structures of various sample distributions associated with a specific class within the latent space presents a challenge using existing visualization methods. This challenge extends to the comparison of multiple sample distributions, limiting the analysis of the latent spaces. Additionally, visually comparing multiple latent spaces side by side can be overwhelming and confusing, especially when dealing with numerous points and varying degrees of distortion [3, 4, 5, 6, 7].
@@
If there is too much information kept, it is hard to understand. At the same time, if the information is filtered enough, some perspective is always lacking. Therefore, some complementary tools might be crucial in interpreting these complex latent spaces.
Figure 2: Overview of the framework to create k* Distribution. We use the learned features of a neural network to compute k* values of individual evaluated samples and then compute the k* distribution for a particular class.
@@
This article introduces a tool designed to enhance understanding of the distribution of samples associated with ‘classes/features’ in the learned latent space generated by neural networks. The proposed approach involves leveraging the local neighborhood relationships among these samples in the latent space. More precisely, the nearest-neighbor method is applied to the learned latent space to identify a k*-nearest neighbor (sample) within the local neighborhood featuring a different class from the test sample. This process evaluates the disruption of the homogeneity within the local neighborhood of the tested sample. Subsequently, a distribution of k* values referred to as the k* distribution is generated, incorporating all the samples belonging to a given class. Through an analysis of the k* distribution, three distinct patterns of distribution of samples in the learned latent space are identified as illustrated in Figure 1:
@@
Pattern A (🟊$\bigstar$) representing Fractured distribution of samples in latent space,
@@
Pattern B ($\clubsuit$) representing Overlapped distribution of samples in latent space, or
@@
Pattern C ($\spadesuit$) representing Clustered distribution of samples in latent space.
@@ II Related Works
II-A Visualizing the Latent Space of Neural Networks using Dimensionality Reduction Techniques:
Given that neural networks operate in a high-dimensional space, visualizing their latent space directly is challenging. Various researchers employ dimensionality reduction techniques to overcome this limitation to represent the latent space in 2 or 3 dimensions. There are a lot of techniques for visualizing the latent space, like, [1, 2, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Many of these algorithms require hyperparameter tuning for optimal visualization, making it challenging to fairly compare latent spaces of varying dimensions. Additionally, while these visualizations effectively capture some perspective on the structure of the learned latent space, using them to draw a clear comparison between different local structures is challenging [3, 4, 5, 6, 7].
The effectiveness of dimensionality reduction techniques becomes apparent in instances when the latent space is well-organized and has successfully assimilated the intended information. In such cases, these visualizations demonstrate utility by aligning with pre-established interpretations. Conversely, when applied to latent space lacking clearly defined pre-established knowledge, the efficacy diminishes for such visualizations where the lack of known structure leads to a ‘blob of points’ in the visualization [7].
II-B Visualizing Association of Features in the Latent Space of Neural Network:
@@
Additionally, interpretable concepts, defined as groups of latent variables in the space that are meaningful, may manifest in the latent space, further contributing to our understanding [31, 32, 36, 37, 38]. In natural language processing, these interpretations often provide insight into learning semantic and linguistic concepts provided by the lexical object [39, 40, 41]. These interpretations are often effectively visualized using analytic systems like [42, 43, 44, 45, 46, 47, 48]; however, the challenge to draw a clear comparison between different local structures and multiple latent spaces persists.
We have seen a few methods significantly contribute to the understanding and visualization of latent spaces using neighborhood-based analyses [49, 50, 51, 52], facilitating the development of more sophisticated models and techniques. They have paved the way for metric learning in neural networks, a domain that focuses on learning distance metrics directly from data [53, 54, 55, 56, 57].
@@
We suggest employing the k* distribution to examine the distribution of samples in the neighborhood of a specific sample and analyze the distribution of samples sharing a predefined label. Our proposed approach presents a framework for effectively comparing latent spaces and sub-spaces with diverse sample distributions (see Figure 2). This framework enables meaningful comparisons between the distribution of samples belonging to different classes and multiple latent spaces, offering additional insights into latent spaces beyond what existing visualizations provide. We hope that our proposed visualization technique using k* distribution can further the advances by providing more insightful analyses of the latent space.
III K* Distribution: Analyzing Homogeneity in the Local Neighborhood of Samples
@@
We propose a methodology based on the local neighborhood to analyze the hyper-dimensional latent space learned by neural networks (see Figure 2). The distribution of samples and clusters in the learned latent space is analyzed by associating them with classes. Through this approach, we gain insights into the distribution patterns within the learned latent space and identify the clusters formed (see Figure 1), thereby enhancing our comprehension of the latent space.
@@
In this context, we introduce the k* distribution by exploring the concept of neighborhoods and their utility in validating the local relationship among features in latent space. The k* distribution is constructed by taking a latent space as input and analyzing the neighborhood of a sample (see Figure 3).
The Nearest Neighbor method, a widely recognized non-parametric technique, enables us to understand the positioning of latent variables in space near a latent variable and their corresponding class distribution. This method aids in gauging the relative distances between latent variables and grouping them into clusters. The underlying principle is grounded in the notion that local neighborhoods offer a reasonable estimate of the sample distribution within the latent space.
@@
The index of the kth nearest neighbor that belongs to a different class than the test sample, disrupting the uniformity of the test sample’s neighborhood, is quantified and termed as the k* value. Essentially, a high k* value indicates that the sample is surrounded by similar points that share the same class. In other words, a high k* value indicates that a neighbor of a different class will be situated a considerable distance from the neighborhood of the specified sample. In particular, the homogeneity of each class cluster within the learned latent space can be assessed using k* (see Figure 3). Additionally, one can ascertain whether a cluster remains cohesive for a given class or if it is fragmented across various spatial locations.
The measurement of the nearest neighbor index allows us to address the sparsity inherent in high-dimensional space. Rather than relying on absolute distance values between points, a relative measure, the neighborhood concept, is utilized. This approach facilitates the comparison of two distinct latent spaces with varying dimensionalities. Importantly, the neighborhood concept is dimensionality-independent, making this technique applicable and effective across low and high dimensions.
@@
Mathematically speaking, let us consider a collection of sample-label pairs $X$: $(x_{1},Y_{1}),(x_{2},Y_{2}),...,(x_{n},Y_{n})$ where $x$ is the input sample, and $Y$ is the label for the input sample. Here, the latent space is embedded with such $x$ points. Let $S$ be the set of all samples $x_{p}\in X$ such that they have the same label $c$, i.e.,
$S_{c}=\{x_{i}\mid\forall x_{i}\in X~{}\text{such that}~{}Y_{i}=c\}.$ (1)
@@
$D(x_{p})=\{\text{distance}(x_{p},x_{i})\mid\forall(x_{i},y_{i})\in X\},$ (2)
where $\text{distance}(a,b)$ is the distance between two samples $a$ and $b$. The commonly employed distance function is the Minkowski distance, a generalization of various distance metrics, including Euclidean and City-block distances. The Minkowski distance of order $r$ between two points $a=(a_{1},a_{2}\ldots a_{d})$ and $b=(b_{1},b_{2}\ldots b_{d})$ in $d$ dimensional space is given by,
$\text{distance}(a,b)_{\text{Minkowski},r}=\left(\sum_{i=1}^{d}|a_{i}-b_{i}|^{r}\right)^{1/r},$ (3)
here, it represents City-block distance ($l_{1}$ norm) when $r=1$; it represents Euclidean distance ($l_{2}$ norm) when $r=2$; and it represents Maximum Norm distance ($l_{\infty}$ norm) when $r=\infty$. $k^{\text{th}}$ neighbour sample $x_{k}^{p}$ to sample $x_{p}$ is defined as,
$x_{k}^{p}=x_{q}\quad\text{where,}\quad q\in\operatorname*{arg\,min}_{x_{q}\in P_{i}}~{}\text{distance}(x_{q},x_{p})$ (4)
such $\text{such that}\quad P_{i}=X-\{x_{j}^{p}\mid\forall j<i\}.$
Similarly, we can define a sorted local neighborhood space $N_{p}$ of sample $(x_{p})$ based on the distance,
@@
$\text{k}^{*}(S_{c})=\left\{\frac{\text{k}^{*}_{p}}{|S_{c}|}\mid\forall x_{p}\in S_{c}\right\},$ (7)
here $|S_{c}|$ is the cardinality of set $S_{c}$ representing the number of samples belonging to class $c$.
Figure 4: Visualization of the distribution of samples in latent space using, (Left) k* distribution, and (Right) Dimensionality Reduction techniques like t-SNE (Top Left), Isomap (Top Right), PCA (Bottom Left), and UMAP (Bottom Right)  of all classes of 16-class-ImageNet for the Logit Layer of ResNet-50 Architecture. Note that the distribution of samples for a particular class is easier to compare using k* distribution than dimensionality reduction techniques.
Based on the defined k* distribution, we can define certain metrics over it as mentioned below,
Mean of k* distribution ($\mu_{k^{*}}$):
@@
Standard Deviation of k* distribution ($\sigma_{k^{*}}$):
@@
Skewness Coefficient of k* distribution ($\gamma_{k^{*}}$):
$\gamma_{k^{*}}=\frac{\frac{1}{|S|}\sum(\text{k}^{*}(S)-\mu_{k^{*}})^{3}}{\left(\frac{1}{|S|}\sum(\text{k}^{*}(S)-\mu_{k^{*}})^{2}\right)^{3/2}}$ (10)
@@
Pattern A (🟊$\bigstar$) Fractured distribution of samples:
@@
Pattern B ($\clubsuit$) Overlapped distribution of samples:
@@
Pattern C ($\spadesuit$) Clustered distribution of samples:
@@
Note: The k* distribution is not a indicator for measuring a neural network’s classification accuracy. The k* distribution visualizes the learned latent space from a local neighborhood perspective, while metrics like accuracy evaluate it from a different perspective. Both multiple well-separated fractured distributions (indication of overclustering) and well-separated homogeneous clusters (indication of optimal clustering) can lead to high classification accuracy.
Limitations: As our approach revolves around identifying surrounding neighbor samples, we inherit the limitations inherent in the nearest-neighbor method. One such trade-off is sacrificing information about the absolute distances between samples and between the distribution of samples.